Circular No. 8019

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ib Further to IAUC 7995, W. M. Wood-Vasey et al. report the discovery

of an apparent supernova (mag 19.8) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Oct. 31.19, 31.21, and 31.23 UT. Confirming NEAT images were taken on Nov. 12 (at mag 18.2), 13 (mag 18.1), and 19 (mag 18.3). The new object is located at α = 0^{h}54^{m}07^{s}.45, δ = +17^{o}33′53′′.7 (equinox 2000.0), which is 1′′ east of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT Palomar images taken on Oct. 3 (limiting mag 21.2 at S/N = 3).

SUPERNOVA 2002ic Wood-Vasey et al. also report the discovery of another apparent su-

pernova (mag 18.5) on unfiltered NEAT images taken on Nov. 13 UT. Confirming NEAT observations were made on Nov. 20 (mag 17.8). SN 2002ic is located at α = 1^{h}30^{m}02^{s}.55, δ = +21^{o}53′06′′.9 (equinox 2000.0), which is 3′′.6 west of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT Haleakala images taken on Oct. 30 (limiting mag 20.0 at S/N = 3).

SUPERNOVA 2002hy IN NGC 3464 S. Benetti, G. Altavilla, A. Pastorello, M. Riello, M. Turatto, and L.

Zampieri, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, on behalf of the ERTN (cf. IAUC 7987), report that a fully reduced CCD spectrum (range 333–1000 nm, resolution 1.2 nm) of SN 2002hy (cf. IAUC 8016), obtained on Nov. 15.33 UT with the European Southern Observatory 3.6-m telescope (+ EFOSC2), shows it to be a peculiar type-Ib supernova caught soon after explosion. The spectrum is dominated by a blue continuum [T_{bb} ∼ 11500 K, once dereddened by E(B −V) ∼ 0.17, as calculated from the equivalent width of the Na I D interstellar band] with overimposed He I lines (at 388.9, 447.1, 501.5, 587.5, 667.8, and 706.5 nm) and their P-Cyg profiles. From the He I 587.5-nm minimum, an expansion velocity of ∼ 9500 km/s is derived. The peaks of the He I emission lines are blueshifted by ∼ 1800 km/s. Narrow absorptions identified with Hα and Hβ are detected at 629.1 and 468.6 nm, respectively, from which an expansion velocity of ∼ 16000 km/s is deduced for Hα.

2002 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green